Question 1

Explain with the help of appropriate schematic diagrams and examples the importance of power factor correction with reference to energy use.

Question 2

What are the different energy use in a classroom?

How can the energy use be reduced? Use appropriate drawings if needed.
Answer #1

Power factor Coxsechoma Power factor u ratis between real system. power and apparent powel in -) A A The closes the voltage a40 kw apparent power 0.95 MSOV P- 40kw ef=0.95 42. KVA tre can see apparent power ů leu than 48 KVA 42. IKVÁ 4%OV: lets consi10 power factor 2015 Now if we introduce a capacitor with same akunk. V 10 XL 4 kune. AKVAR S= Tiot (9-9) Pf = 1%0 Now let uV Xc Cuso) 36,864 VAR - Xe = сос whe 27X 60 X 6.25 C=424.4uF power i mos Pincreases. As the energy So if we place Capacitor C